# Privacy risk in GeoData: A survey

### v1

@@ 3.3.1. Data sets
GeoPriv (Pablo Duque and Pugliese, 2023)
A QGIS 3+ [^1] https://qgis.org/en/site/ plugin that offers three location privacy methods (i.e., Spatial Clustering (Zurbaran and Wightman, 2017), NRandK (Zurbarán et al., 2018), and Laplace Noise (or Geo indistinguishability) (Andrés et al., 2013)). These methods are mainly designed for location-based services and focus on Vector Layers with Point geometries, utilising obfuscation and k-anonymity.
https://github.com/Diuke/GeoPriv
MaskMy.XYZ (Swanlund et al., 2020a)
A web application designed for easy geographic masking. This tool uses donut geomasking and includes metrics for privacy protection and information loss, enabling users to adjust masking parameters iteratively based on these metrics.
https://maskmy.xyz/
PrivyTo (McKenzie et al., 2022)
This tool is designed to protect privacy when sharing personal location data. Users can generate multiple obfuscated views using various geomasking techniques (i.e., random perturbations), tailoring each for specific recipients. The application utilise encryption for secure data transfer.
https://privyto.me/site
Privy (Ajayakumar et al., 2019)
A geomasking tool mainly designed for healthcare. It transfers key location information to a different geographic location, analysing the data, and finally returns the results to the original location. Privy relocates coordinates using a combination of random translations and rotations while maintaining the connectivity of spatial point data.
https://github.com/ghhlab/confidentiality
MapSafe (Sharma et al., 2023)
A web-based application to share encrypted geospatial information that uses donut masking and hexagonal binning for dataset obfuscation, followed by a multi-level encryption scheme. This tool enables data owners to share data at their preferred level of detail.
https://www.mapsafe.xyz/
StppSim (Adepeju, 2023)
A tool designed for synthesising fine-grained spatio-temporal crime records. This tool allows to generate synthetic data from a sample dataset or based on user-provided spatio-temporal characteristics.
CRAN (https://cran.r-project.org/web/packages/stppSim/index.html) and GitHub (https://github.com/cran/stppSim)
GeoMasker (Chen et al., 2017)
A geospatial tool designed to balance the protection of geoprivacy and the accuracy of spatial patterns within a GIS. The tool employs carefully calibrated geomasking parameters, including grid size (GS), estimated K-anonymity (K), and D statistics, to mask the residential locations within GIS software. GeoMasker is valuable for health analytics enabling secure information processing and sharing while preserving privacy and spatial patterns.
http://idv.sinica.edu.tw/tachien/geomasker
3.3.2. Implementations
This dimension specifies the applications and prototypes that have been developed and implemented for various geoprivacy techniques in order to conduct experimental evaluation. Some solutions proposed in the literature provide only algorithmic details but they have not been evaluated experimentally, or no details about their implementation have been published. Table 3 provides a brief description of existing tools that implemented with different geomasking techniques.
@@ 3.3.3. Application Domains
Geospatial information, spanning a diverse range of location-based data, plays a pivotal role across various application domains, including but not limited to disaster management, health analytics, urban planning and management, and agriculture. These distinct realms are explored in greater detail in the subsequent subsections.
@@ 3.3.3.4 Agriculture
Geospatial technology plays a pivotal role in agriculture. Through the utilisation of remotely sensed images acquired from satellites, aerial platforms, and GPS-tagged drones, geospatial technology facilitates the mapping and monitoring of various agricultural parameters, including soil moisture, nutrient availability, crop types, growth stages, and disease stress. The integration of geospatial technology into precision agriculture not only supports sustainable development goals, but also aids in conserving biodiversity, mitigating environmental impacts, and promoting sustainable land management practices worldwide (Pandey and Pandey, 2023; Obi Reddy et al., 2023).
@@
4. A survey of geomasking techniques
In this section, a comprehensive survey of recent geomasking techniques is conducted, examining the methods employed to preserve geospatial privacy while maintaining data utility and accuracy. We provide the details of the existing literature following the categorisation we used for different geomasking techniques on Section 3.1.
@@ 4.1. Random perturbation
Circular masking
with fixed radius
@@
Circular masking
with random radius
@@
Random directions
within annual
(Donut masking)
@@
Spatially adaptive random
perturbation (SARP)
@@
Weighted random
perturbation (WRP)
@@
Bimodal Gaussian
displacement
@@
Density based Gaussian
spatial skew
@@
Multiple risk factor
adaptive displacement
@@
Table 4 outlines different papers that proposed or utilised a random perturbation technique. Kwan et al. (Kwan et al., 2004a) investigated the use of circular masking with a fixed and random radius $r$. According to the investigation a larger $r$ for the same mask leads to more deviation from the pattern obtained from the unmasked data. Further, the study suggests that circular and random radius masks, in general, yield better results than the weighted masks with the same $r$ (Allshouse et al., 2010).
@@
Stinchcomb (Stinchcomb, 2004) was the first to explore donut masking for displacing points in random directions within an annulus. However, in cases where population density is not evenly distributed, donut masking could lead to reveal information about individuals. To mitigate this, Allshouse et al. (Allshouse et al., 2010) suggested tripling the minimum displacement area within the donut for heterogeneous populations, ensuring privacy preservation with less than a 1% error rate.
@@
Further, Hampton et al. (Hampton et al., 2010) examines the donut method for disease mapping and compares its performance with other geomasking techniques (i.e., aggregation and random perturbation) in terms of cluster detection and patient geoprivacy protection. Their investigation shows that the donut method and random perturbation outperformed aggregation in measuring hot spot/clustered areas. However, the donut method provides a higher level of privacy preservation compared to the random perturbation, with only a slight reduction in cluster detection performance, especially in areas where individual privacy is at a higher risk.
@@
Lu et al. (Lu et al., 2012) explored a masking technique based on Spatially adaptive random perturbation (SARP). As we discussed in Section 3.1, the SARP technique defines the perturbation zone based on the actual distribution of residential addresses (risk location) instead of people (risk population). It employs ”donut-shaped” spatially varied perturbation zones rather than ”pancake-shaped” which allows a point to be moved to a nearby location within its immediate surroundings. The study findings shows that SARP geomasking, utilising actual street addresses, offers better privacy protection compared to geomasking based on population size. While the SARP techniques do not significantly alter clustering patterns on a global scale, the geomasked data tends to exhibit greater clustering compared to the real distribution.
@@
Cassa et al. (Cassa et al., 2006), Zandbergen (Zandbergen, 2014), and Cassa et al.(Cassa et al., 2008) investigated the masking methods based on Gaussian displacement. However, as these studies highlighted the Gaussian displacement’s main challenge lies in the limited point displacement due to its concentration around the mean. As highlighted by Zandbergen, donut masking and bimodal Gaussian displacement are very similar in terms of the general area where masked locations are placed relative to the original locations (Zandbergen, 2014).
Murad et al. (Murad et al., 2014) investigated the use of triangular displacement as a masking technique for geodata. The authors suggested a method that evaluates various risk factors to determine the appropriate displacement distances for individual location points. It begins by assessing the re-identification risk of each point, utilising a risk factor of 1:20. If an area has fewer than 20 nearby points within a one-mile radius, it is categorised as ”cannot be de-identified,” signifying a higher re-identification probability.
The method also considers the sensitivity level of the data, adding specific displacement distances for low and high sensitivity information (300m and 500m, respectively). Additionally, it takes into account the research type, distinguishing between advanced and traditional research. In cases of advanced research, an extra 100m is added to the displacement distance, reflecting the potentially greater complexity and nuanced nature of the findings. Quasi-identifiers, which pose a combined risk when aggregated, prompt the algorithm to add 100m to the displacement distance, while grouping these quasi-identifiers until at least 20 individuals are in each group.
@@
To further minimise external risks, an extra 100m is added to the displacement distance. The algorithm then considers the intended release of the dataset, differentiating between public and restricted access. For public releases, an additional 100m is included in the displacement distance. The resulting displacement distance, treated as a range with lower and upper limits ($R_{Lower}$, $R_{Upper}$), is employed in the Triangular Displacement method. This method, based on the Pythagorean Equation ($x^{2}+y^{2}=r^{2}$), introduces a random yet controlled shift in the original location points, providing a versatile and context-aware approach to spatial data masking.
@@ 4.2. Masking with pre-defined locations
As we described in Section 3.1, techniques that mask with pre-defined location move a given point to a new location, which is randomly chosen from all the possible locations of a similar type that exist within a specific region (such as a circle or an annulus) surrounding the original location. Zhang et al. (Zhang et al., 2017) was the first to use location swapping as a masking technique for geospatial data in the health analytic domain. The location swapping methods provide a more realistic scenario in terms of the displacement location selection for masking. Because location swapping provides an additional filter to ensure masked locations fall within predefined areas of interest.
However, as the authors highlighted, if an attacker able to determine the buffer distance used to relocate a point, it might be possible to re-identify the original locations. The same authors also explored the location swapping with donut masking to set a minimum distance preventing points from being swapped with any locations within this specified range. The authors employed an internal buffer size half that of the external buffer.
Swanlund et al. (Swanlund et al., 2020b) introduced street masking which uses a road network as a basis for masking. The authors used OpenStreetMap road network to replace the original location with coordinates sampled from a set within a predefined radius. As the study suggests, street masking performs similarly to population-based donut geomasking with regard to privacy protection, achieving comparable k-anonymity values at similar median displacement distances.
@@ 4.3. Affine transformation
As we discussed in Section 3.1, affine transformations can be global, applied uniformly to all points, or local, affecting a smaller area, based on the chosen scale of implementation. These techniques preserve the spatial structure of the data, proving useful for spatial analysis and visualisation methods such as clustering, while maintaining the number of records and avoiding random relocation of data (Ajayakumar et al., 2019). However, in these techniques, if the transformation matrix is known, retrieving the coordinates of the original records from the perturbed dataset is simple. Therefore, affine transformations are not a widely used methods (Wang et al., 2022).
Recently, Wang et al. (Wang et al., 2022) conducted a study that investigated the use of translation, rotating, and scaling techniques for geomasking. The authors highlighted that the coordinates of the original records can be easily recovered from the geomasked dataset if the attacker knows the transformation matrix and even the identification of only a few points could make it possible to re-identify the entire point pattern.
Leitner and Curtis were the first to explore flipping as a geomasking technique (Leitner and Curtis, 2004). The authors stated that most variants (e.g. flipping horizontally and flipping vertically) of the flipping method are designed for publishing maps and not for releasing the coordinates (Leitner and Curtis, 2006). Broen et al. (Broen et al., 2021) investigated the impact of horizontal shear as a masking technique. According to the authors, horizontal shear conferred the greatest anonymity using the k-anonymity metric and maintained some spatial patterns compared to other affine transformation techniques. However, the method is not secure, as points can be trivially re-identified if the angle of the shearing can be determined.
@@ 4.4. Spatial aggregation and Clustering
Spatial aggregation and clustering techniques are more popular than other masking techniques due to the aggregation of aggregates individual location points within specific geographic units to ensure anonymity and prevent the disclosure of exact locations. Table 5 outlines the literature that uses spatial aggregation and clustering techniques for masking geodata.
@@
Adaptive areal
elimination (AAE)
@@
Adaptive areal
masking (AAM)
@@
Adaptive voronoi
masking (AVM)
@@
As explored by (Leitner and Curtis, 2004; Gupta and Rao, 2020), grid masking disparate every original location data point to uniform grid cells. However, grid masking disrupt the original data points pattern and reduce the data utility for further research and analysis. Lagonigro et al. (Lagonigro et al., 2017) proposed a methodology based on quadtree hierarchical geographic data structures to create a varying size grid adapted to local area densities. As a hierarchical grid system, it can easily be aggregated into a fixed size grid on a lower resolution level. The methodology has been implemented using the R software and applied to the Catalonian 2014 population register data. Clarke (Clarke, 2016) investigated a multiscale geomasking technique by which locations are converted to Military Grid Reference System (MGRS) coordinates provides a unique amount of control over the adjusted locations. Despite the computing time required for processing, the method offers new possibilities for protecting sensitive geospatial data, overcoming disadvantages of geometric point masking methods.
@@
Voronoi masking (VM) is a spatial tessellation technique which can retain the spatial integrity of the original data. Seidl et al. (Seidl et al., 2015) was the first to introduce the idea of VM. One advantage of this method is that points in neighbouring polygons are displaced to the same position, enhancing their K-anonymity. However, VM can dislocate some data points to large distances if the area contain scatted positions, which will change spatial patterns. This mask was automated in Python for ArcGIS Pro and ArcMap.
@@
Kounadi and Leitner (Kounadi and Leitner, 2016) proposed Adaptive areal elimination (AAE) which creates areas of a minimum K-anonymity and then original points are either randomly perturbed within the areas or aggregated to the median centers of the areas. According to the authors, AAE introduces less spatial error than the donut mask for all the measures of spatial error. Polzin (Polzin, 2020) integrated the principles of AAE and VM to introduce a masking technique called Adaptive voronoi masking (AVM). AVM addresses topographical errors by relocating data points to the nearest street intersection, minimising false re-identification and illogical placements in water bodies or forests.
@@
Gupta and Rao (Gupta and Rao, 2020) proposed a method very similar to VM called RDV masking. Depending on the region size, for large regions, points are moved to the nearest segment of the edges of the polygon. Charleux and Schofield (Charleux and Schofield, 2020) introduced a variation of AAE called Adaptive Areal Masking (AAM). This method tries to merge polygons not by the largest shared border but with the polygon with the closest centroid. However, in configurations where the number of points to anonymise is high enough and the number of population polygons is low enough, AAE can outperform AAM. Similar to AAE, anonymisation polygons generated by the AAM method should not be disclosed to end users to minimise any potential reidentification attacks.
@@
Ajayakumar et al. (Ajayakumar et al., 2023) proposed to use Zip4 codes as a spatial aggregation unit for sharing fine-scale health data. Analysing 1.6 million Ohio Zip4 codes, the research reveals Zip4’s high spatial precision, particularly in urban locales, making it well-suited for detailed spatial analysis than common aggregation units. However, the study underscores a potential privacy risk, noting that Zip4 centroids have a higher likelihood of compromising spatial anonymity, with 73% of addresses exhibiting a low spatial k-anonymity value.
@@
Differential private (DP) clustering focuses on the critical challenge of clustering sensitive data while preserving individual privacy, which holds significant importance in machine learning and data analysis applications. Su et al. (Su et al., 2017) delves into the limitations of current state-of-the-art approaches for k-means clustering, particularly focusing on methodologies that employ a single-workload technique. (Ghazi et al., 2020) delves into differentially private clustering, focusing on fundamental clustering problems such as Euclidean DensestBall, 1-Cluster, k-means, and k-median. (Lu and Shen, 2020; Cohen et al., 2021) proposes a new approach to integrating differential privacy into clustering models, addressing the limitations of existing differentially private clustering algorithms. (Yang et al., 2022) addresses the emerging field of privacy-preserving data analysis, specifically focusing on constructing a privacy-preserving K-means clustering protocol that ensures local privacy for user data.
@@
Houfaf-Khoufaf et al. (Houfaf-Khoufaf et al., 2021) studied the combination of aggregation techniques with perturbation mechanisms. The study also introduces the simulated cluster crowding (SCC) method, a novel approach inspired by the simulated crowding of bike GPS tracks. SCC generates new random points within existing clusters’ extents to achieve both high k-anonymity and cluster preservation. The clusters are initially generated using the DBSCAN algorithm, distinguishing between outlier points and those within clusters. Outlier points are masked using the bimodal Gaussian perturbation method, while points within clusters are masked by generating the same number of new points randomly inside the polygon extent of each cluster. Three versions of SCC are proposed: SCCv1 generates points uniformly within the cluster area, SCCv2 introduces buffer areas around initial points to ensure a minimum distance, and SCCv3 dilates roads and intersects them with clusters to generate fake points. The methods are evaluated based on k-anonymity, cluster preservation, and privacy guarantee, demonstrating their effectiveness in maintaining privacy while preserving the spatial distribution of clusters.

### v2

@@ 3.3.2. Implementations
GeoPriv (Pablo Duque and Pugliese, 2023) A QGIS 3+ [^1] https://qgis.org/en/site/ plugin that offers three location privacy methods (i.e., Spatial Clustering (Zurbaran and Wightman, 2017), NRandK (Zurbarán et al., 2018), and Laplace Noise (or Geo indistinguishability) (Andrés et al., 2013)). These methods are mainly designed for location-based services and focus on Vector Layers with Point geometries, utilising obfuscation and k-anonymity. https://github.com/Diuke/GeoPriv
MaskMy.XYZ (Swanlund et al., 2020a) A web application designed for easy geographic masking. This tool uses donut geomasking and includes metrics for privacy protection and information loss, enabling users to adjust masking parameters iteratively based on these metrics. https://maskmy.xyz/
PrivyTo (McKenzie et al., 2022) This tool is designed to protect privacy when sharing personal location data. Users can generate multiple obfuscated views using various geomasking techniques (i.e., random perturbations), tailoring each for specific recipients. The application utilises encryption for secure data transfer. https://privyto.me/site
Privy (Ajayakumar et al., 2019) A geomasking tool mainly designed for healthcare. It transfers key location information to a different geographic location, analysing the data, and finally returns the results to the original location. Privy relocates coordinates using a combination of random translations and rotations while maintaining the connectivity of spatial point data. https://github.com/ghhlab/confidentiality
MapSafe (Sharma et al., 2023) A web-based application to share encrypted geospatial information that uses donut masking and hexagonal binning for dataset obfuscation, followed by a multi-level encryption scheme. This tool enables data owners to share data at their preferred level of detail. https://www.mapsafe.xyz/
StppSim (Adepeju, 2023) A tool designed for synthesising fine-grained spatiotemporal crime records. This tool allows to generate synthetic data from a sample dataset or based on user-provided spatiotemporal characteristics. https://github.com/cran/stppSim
GeoMasker (Chen et al., 2017) A geospatial tool designed to balance the protection of geoprivacy and the accuracy of spatial patterns within a GIS. The tool employs carefully calibrated geomasking parameters, including grid size (GS), estimated K-anonymity (K), and D statistics, to mask the residential locations within GIS software. GeoMasker is valuable for health analytics, enabling secure information processing and sharing while preserving privacy and spatial patterns. http://idv.sinica.edu.tw/tachien/geomasker
@@ 3.3.3. Application Domains
Geospatial information, spanning a diverse range of location-based data, plays a pivotal role across various application domains, including but not limited to disaster management, health analytics, urban planning and management, and agriculture. These distinct sectors are explored in greater detail in the subsequent subsections.
@@ 3.3.3.4 Agriculture
Geospatial technology plays a pivotal role in agriculture. Through the utilisation of remotely sensed images acquired from satellites, aerial platforms, and GPS-tagged drones, geospatial technology facilitates the mapping and monitoring of various agricultural parameters, including soil moisture, nutrient availability, crop types, growth stages, and disease stress. The integration of geospatial technology into precision agriculture not only supports sustainable development goals but also aids in conserving biodiversity, mitigating environmental impacts, and promoting sustainable land management practices worldwide (Pandey and Pandey, 2023; Obi et al., 2023).
@@
4. Existing literature on geomasking techniques
In this section, a comprehensive survey of recent geomasking techniques is conducted, examining the methods employed to preserve geospatial privacy while maintaining data utility and accuracy. We provide the details of the existing literature following the categorisation in Section 3.1.
@@ 4.1. Random perturbation
Circular masking with fixed radius
@@
Circular masking with random radius
@@
Random directions within annual (Donut masking)
@@
Spatially adaptive random perturbation (SARP)
@@
Weighted random perturbation (WRP)
@@
Bimodal Gaussian displacement
@@
Density based Gaussian spatial skew
@@
Multiple risk factor adaptive displacement
@@
Table 4 outlines different papers that proposed or utilised a random perturbation technique. Kwan et al. (Kwan et al., 2004a) investigate the use of circular masking with a fixed and random radius $r$. According to the investigation, a larger $r$ for the same mask leads to more deviation from the pattern obtained from the unmasked data. Further, the study suggests that circular and random radius masks, in general, yield better results than the weighted masks with the same $r$ (Allshouse et al., 2010).
@@
Stinchcomb (Stinchcomb, 2004) was the first to explore donut masking for displacing points in random directions within an annulus. However, in cases where population density is not evenly distributed, donut masking could lead to revealing information about individuals. To mitigate this, Allshouse et al. (Allshouse et al., 2010) suggest tripling the minimum displacement area within the donut for heterogeneous populations, ensuring privacy preservation with less than a 1% error rate.
@@
Further, Hampton et al. (Hampton et al., 2010) examine the donut method for disease mapping and compares its performance with other geomasking techniques (i.e., aggregation and random perturbation) in terms of cluster detection and patient geoprivacy protection. Their investigation shows that the donut method and random perturbation outperformed aggregation in measuring hot spot/clustered areas. However, the donut method provides a higher level of privacy preservation compared to the random perturbation, with only a slight reduction in cluster detection performance, especially in areas where individual privacy is at a higher risk.
@@
Lu et al. (Lu et al., 2012) explore a masking technique based on Spatially adaptive random perturbation (SARP). As we have discussed in Section 3.1, the SARP technique defines the perturbation zone based on the actual distribution of residential addresses (risk location) instead of people (risk population). It employs ”donut-shaped” spatially varied perturbation zones rather than ”pancake-shaped” which allows a point to be moved to a nearby location within its immediate surroundings. The study findings show that SARP geomasking, utilising actual street addresses, offers better privacy protection compared to geomasking based on population size. While the SARP techniques do not significantly alter clustering patterns on a global scale, the geomasked data tends to exhibit greater clustering compared to the real distribution.
@@
Cassa et al. (Cassa et al., 2006), Zandbergen (Zandbergen, 2014), and Cassa et al.(Cassa et al., 2008) investigate the masking methods based on Gaussian displacement. However, these studies highlight that the main challenge of Gaussian displacement lies in the limited point displacement due to its concentration around the mean. As highlighted by(Zandbergen, 2014), donut masking and bimodal Gaussian displacement are very similar in terms of the general area where masked locations are placed relative to the original locations.
@@
Murad et al. (Murad et al., 2014) investigate the use of triangular displacement as a masking technique for geodata. The authors suggest a method that evaluates various risk factors to determine the appropriate displacement distances for individual location points. The method considers the sensitivity level of the data, adding specific displacement distances for low and high sensitivity information. Quasi-identifiers, which pose a combined risk when aggregated, prompt the algorithm to add additional distance to the displacement distance, while grouping these quasi-identifiers until a preset number of individuals are in each group. The algorithm then considers the intended release of the dataset, differentiating between public and restricted access. The resulting displacement distance, treated as a range with lower and upper limits ($R_{Lower}$, $R_{Upper}$), is employed in the Triangular Displacement method. This method, based on the Pythagorean Equation, introduces a random yet controlled shift in the original location points, providing a versatile and context-aware approach to spatial data masking.
@@ 4.2. Masking with pre-defined locations
Zhang et al. (Zhang et al., 2017) was the first to use location swapping as a masking technique for geospatial data in the health analytic domain. This technique offers a more realistic scenario for displacement location selection. However, as the authors highlighted, if an attacker is able to determine the buffer distance used to relocate a point, it might be possible to re-identify the original locations. The same authors also explored the location swapping with donut masking to set a minimum distance preventing points from being swapped with any locations within this specified range. The authors employed an internal buffer size half that of the external buffer.
Swanlund et al. (Swanlund et al., 2020b) introduce street masking, which uses a road network as a basis for masking. The authors use OpenStreetMap road network to replace the original location with coordinates sampled from a set within a predefined radius. As the study suggests, street masking performs similarly to population-based donut geomasking with regard to privacy protection, achieving comparable k-anonymity values at similar median displacement distances.
Geo-indistinguishability, as proposed by Andrés et al. (Andrés et al., 2013), is a privacy notion proposed for preserving privacy in location-based services. It offers stronger protection than spatial k-anonymity by safeguarding against adversaries with any current or future side information. The notion is formalised as a user enjoying $\epsilon r$-privacy within a radius $r$, where $\epsilon$ determines the level of privacy per unit distance, with smaller values of $\epsilon$ providing higher privacy protection. It achieves privacy protection by relocating individual point locations, ensuring that within a certain radius, any two points yield similar locations after relocation.
Lin (Lin, 2023) introduces geo-indistinguishable masking as a robust solution to safeguard privacy in spatial point mapping. By leveraging the concept of geo-indistinguishability, the approach ensures strong privacy protection while preserving spatial point patterns through a two-step process. Initially, masking areas are generated around each spatial point, comprising a set of candidate locations for potential relocation. Subsequently, an optimisation model is employed to strategically select the optimal candidate location within each masking area, minimising displacement distance while adhering to the stringent privacy criterion of geo-indistinguishability. While sharing similarities with location swapping methods, which also relocate spatial points within masking areas, geo-indistinguishable masking distinguishes itself by prioritising optimal candidate selection over random assignment, thereby offering superior privacy guarantees. Despite its computational complexity and potential limitations regarding temporally correlated spatial points and risk of false identification, the method represents a significant advancement in the realm of privacy-preserving spatial analysis, essential for upholding individual privacy rights in an era of burgeoning geospatial data utilisation.
@@ 4.3. Affine transformation
Recently, Wang et al. (Wang et al., 2022) conduct a study that investigates the use of translation, rotating, and scaling techniques for geomasking. The authors highlight that the coordinates of the original records can be easily recovered from the geomasked dataset if the attacker knows the transformation matrix, and even the identification of only a few points could make it possible to re-identify the entire point pattern.
Leitner and Curtis were the first to explore flipping as a geomasking technique (Leitner and Curtis, 2004). The authors state that most variants (e.g. flipping horizontally and flipping vertically) of the flipping method are designed for publishing maps and not for releasing the coordinates (Leitner and Curtis, 2006). Broen et al. (Broen et al., 2021) investigate the impact of horizontal shear as a masking technique. According to the authors, horizontal shear confers the greatest anonymity using the k-anonymity metric and maintains some spatial patterns compared to other affine transformation techniques. However, the method is not secure, as points can be trivially re-identified if the angle of the shearing can be determined.
@@ 4.4. Spatial aggregation and Clustering
Table 5 outlines the literature that uses spatial aggregation and clustering techniques for masking geodata.
@@
Adaptive areal elimination (AAE)
@@
Adaptive areal masking (AAM)
@@
Adaptive voronoi masking (AVM)
@@
As explored by (Leitner and Curtis, 2004; Gupta and Rao, 2020), grid masking disparate every original location data point to uniform grid cells. However, grid masking disrupts the original data points’ pattern and reduces the data utility for further research and analysis. Lagonigro et al. (Lagonigro et al., 2017) propose a methodology based on quadtree hierarchical geographic data structures to create a varying-size grid adapted to local area densities. As a hierarchical grid system, it can easily be aggregated into a fixed-size grid on a lower resolution level. The methodology has been implemented using the R software and applied to the Catalonian 2014 population register data. Clarke (Clarke, 2016) investigates a multiscale geomasking technique, whereby locations are converted to Military Grid Reference System (MGRS) coordinates, providing a unique level of control over the adjusted locations. Despite the computing time required for processing, the method offers new possibilities for protecting sensitive geospatial data, overcoming disadvantages of geometric point masking methods.
@@
Voronoi masking (VM) is a spatial tessellation technique that can retain the spatial integrity of the original data. Seidl et al. (Seidl et al., 2015) was the first to introduce the idea of VM. One advantage of this method is that points in neighbouring polygons are displaced to the same position, enhancing their K-anonymity. However, VM can dislocate some data points to large distances if the area contains scatted positions, which will change spatial patterns. This mask was automated in Python for ArcGIS Pro and ArcMap.
@@
Kounadi and Leitner (Kounadi and Leitner, 2016) propose Adaptive areal elimination (AAE) which creates areas of a minimum K-anonymity, and then original points are either randomly perturbed within the areas or aggregated to the median centers of the areas. According to the authors, AAE introduces less spatial error than the donut mask for all the measures of spatial error. Polzin (Polzin, 2020) integrates the principles of AAE and VM to introduce a masking technique called Adaptive voronoi masking (AVM). AVM addresses topographical errors by relocating data points to the nearest street intersection, minimising false re-identification and illogical placements in water bodies or forests.
@@
Gupta and Rao (Gupta and Rao, 2020) propose a method very similar to VM called RDV masking. Depending on the region size, for large regions, points are moved to the nearest segment of the edges of the polygon. Charleux and Schofield (Charleux and Schofield, 2020) introduce a variation of AAE called Adaptive Areal Masking (AAM). This method tries to merge polygons not by the largest shared border but with the polygon with the closest centroid. However, in configurations where the number of points to anonymise is high enough and the number of population polygons is low enough, AAE can outperform AAM. Similar to AAE, anonymisation polygons generated by the AAM method should not be disclosed to end users to minimise any potential reidentification attacks.
@@
Ajayakumar et al. (Ajayakumar et al., 2023) propose to use Zip4 codes as a spatial aggregation unit for sharing fine-scale health data. Analysing 1.6 million Ohio Zip4 codes, the research reveals Zip4’s high spatial precision, particularly in urban locales, making it well-suited for detailed spatial analysis than common aggregation units. However, the study underscores a potential privacy risk, noting that Zip4 centroids have a higher likelihood of compromising spatial anonymity, with 73% of addresses exhibiting a low spatial k-anonymity value.
@@
Differential private (DP) clustering focuses on the critical challenge of clustering sensitive data while preserving individual privacy, which holds significant importance in machine learning and data analysis applications. Su et al. (Su et al., 2017) delve into the limitations of current state-of-the-art approaches for k-means clustering, particularly focusing on methodologies that employ a single-workload technique. Ghazi et al. (Ghazi et al., 2020) delve into differentially private clustering, focusing on fundamental clustering problems such as Euclidean DensestBall, 1-Cluster, k-means, and k-median. Cohen et al. (Cohen et al., 2021) and Lu and Shen (Lu and Shen, 2020) propose a new approach to integrating differential privacy into clustering models, addressing the limitations of existing differentially private clustering algorithms. Yang et al. (Yang et al., 2022) address the emerging field of privacy-preserving data analysis, specifically focusing on constructing a privacy-preserving K-means clustering protocol that ensures local privacy for user data.
@@
Houfaf-Khoufaf et al. (Houfaf-Khoufaf et al., 2021) study the combination of aggregation techniques with perturbation mechanisms. The study also introduces the simulated cluster crowding (SCC) method, a novel approach inspired by the simulated crowding of bike GPS tracks. SCC generates new random points within existing clusters’ extents to achieve both high k-anonymity and cluster preservation. The clusters are initially generated using the DBSCAN algorithm, distinguishing between outlier points and those within clusters. Outlier points are masked using the bimodal Gaussian perturbation method, while points within clusters are masked by generating the same number of new points randomly inside the polygon extent of each cluster. Three versions of SCC are proposed: SCCv1 generates points uniformly within the cluster area, SCCv2 introduces buffer areas around initial points to ensure a minimum distance, and SCCv3 dilates roads and intersects them with clusters to generate fake points. The methods are evaluated based on k-anonymity, cluster preservation, and privacy guarantee, demonstrating their effectiveness in maintaining privacy while preserving the spatial distribution of clusters.